OpenStudy (anonymous):

all solutions for 2cos3theta=1

OpenStudy (mertsj):

Divide both sides by 2.

OpenStudy (anonymous):

what about after that???

OpenStudy (mertsj):

Do you know an angle whose cosine is 1/2?

OpenStudy (anonymous):

60 and 300??

OpenStudy (anonymous):

what do you do with the 3 theta??

OpenStudy (mertsj):

So we know that 3 theta is 60 and 300 but... if you want all of the solutions of theta that are between 0 and 360 you must find all the 3 theta solutions between 0 and 1080.

OpenStudy (anonymous):

???

OpenStudy (mertsj):

\[3\theta =60^{o},300^{o},420^{o},660^{o}, 780^{o}, 1020^{o}\]

OpenStudy (anonymous):

THANK YOU SO MUCH :) what about tan3\[\tan3\theta=\sqrt{3}/3\]

OpenStudy (mertsj):

\[\theta = 20^{o}, 100^{o},140^{o}, 220^{o}, 260^{o}, 340^{o}+2\pi n\]

OpenStudy (mertsj):

\[3\theta = 30^{o},150^{o},390^{o},510^{o},750^{o},870^{o}\]

OpenStudy (anonymous):

what about\[2\tan^2\theta+\tan \theta=2-2\tan\] using the quadratic formula please show work thank you so much!!!!

OpenStudy (mertsj):

\[\theta = 10^{o}, 50^{o}, 130^{o},170^{o},250^{o}, 290^{o}+2\pi n\]

OpenStudy (mertsj):

Get it equal to 0.

OpenStudy (anonymous):

okay thanks!!!!

OpenStudy (mertsj):

yw